1.The turkey industry is perfectly competitive. Gobbler’s Turkey Farm is one of 1,500 turkey farms in the industry. Each firm has the same cost curves. The market is currently in long-run equilibrium with a market price of $6 per pound and a total market production of 12,000,000 pounds.

a. How many pounds of turkey is Gobbler’s Turkey Farm currently producing?

b.What is Gobbler’s Turkey Farm’s ATC at this output level?

c.What is the Economic Profit of Gobbler’s Turkey Farm?

d.Draw two graphs: 1) Gobbler’s Turkey Farm MR, MC, and ATC curves and 2) the market supply and demand curves for turkey.

Be sure to correctly label all axes and indicate equilibrium.

2.Suppose the American Medical Association (AMA) reports that consumption of turkey reduces the chances of a heart attack.

Indicate on your graphs the expected short-run effects from this type of news on the market for turkeys and Gobbler’s Turkey Farm.

3. Given your answer in 2, what do you expect to happen in the long-run in the turkey market?

a. Show on your graphs the expected changes that will occur?

b. What will be the long-run market price in the turkey market?

c.What is the Economic Profit of Gobbler’s Turkey Farm?
Answer #1

Answer to Question No. 1

Pounds of turkey is Gobbler’s Turkey Farm currently producing is 12000000/1500 = 8000 pounds (Total output/number of firms)

Firm's ATC is equal to the price of the product when the firm is in long run equilibrium. So ATC=P =$6.

Economic Profits in a situation when firm is in long run equilibrium is ZERO. The formula for profits = (P-ATC)*Q = (6-6)*8000 = 0